# GenExam: A Multidisciplinary Text-to-Image Exam

### v2

@@ 4.2 Main Results
Semantic
Correctness
Spelling
Logical
Consistency
Readability
@@
Relaxed Score 0.6746 <<0.0001 0.8217 <<0.0001 0.8444 <<0.0001
@@
Table 3: Mean absolute error (MAE) and correlations between human and automatic evaluation. Our evaluation shows strong alignment with human preferences, evidented by low MAE and high correlations.
@@ 4.3 Human Alignment
We conduct a human alignment experiment to examine the validity of MLLM-as-a-judge, following the paradigms in  Zhao et al. (2025); Sun et al. (2025). Five experts are asked to score 250 randomly sampled images generated by GPT-Image-1. They first provide an overall rating between 1 and 10, and then follow the automatic evaluation protocol to annotate scoring points, spelling, logical consistency and readability. The average scores among all experts are used to calculate the mean absolute error (MAE) and correlations between MLLM-judge predicted scores and human scores. In Tab. 3(a), the results show a low MAE across the four dimensions, indicating high precision of our evaluation pipeline.
@@
We then use the overall rating to calculate correlations (Kendall’s $\tau$, Spearman’s $\rho$ and Pearson’s $r$) between human scores and four automatic metrics: our relaxed score, our semantic correctness, VQA score Lin et al. (2024), and CLIP score Hessel et al. (2021). In Tab. 3(b), our relaxed score achieves high correlations with human judgments with $p<0.05$, slightly better than using solely semantic correctness without visual plausibility, while CLIP score and VQA score fail to capture the correctness of multidisciplinary images.
4.4 Ablation on the Evaluator Model
Evaluator Model
GT Image
Input
Scoring Points
Accuracy $\uparrow$
Plausibility
MAE $\downarrow$
Kendall
$\tau\uparrow$
Spearman
$\rho\uparrow$
Pearson
$r\uparrow$
API Cost
(\$/img)
Time
(s/img)
Closed-source MLLMs
GPT-5 + minimal reasoning effort 84.2 0.28 0.6066 0.7521 0.7860 0.0144 35
GPT-5 + low reasoning effort 88.5 0.20 0.6746 0.8217 0.8444 0.0242 72
GPT-5 + low reasoning effort $\times$ 86.2 0.23 0.6141 0.7753 0.7976 0.0231 59
GPT-5 + medium reasoning effort 86.7 0.21 0.6617 0.8224 0.8358 0.0389 101
GPT-5 + high reasoning effort 87.6 0.20 0.6647 0.8190 0.8304 0.0654 186
GPT-5-mini + minimal reasoning effort 79.0 0.35 0.5446 0.6996 0.7185 0.0039 11
GPT-5-mini + low reasoning effort 84.3 0.29 0.6002 0.7729 0.7806 0.0052 19
GPT-5-mini + medium reasoning effort 84.2 0.25 0.6313 0.8025 0.8165 0.0079 36
GPT-5-mini + high reasoning effort 83.4 0.26 0.5906 0.7611 0.7811 0.0196 119
o3 82.0 0.23 0.6082 0.7706 0.7895 0.0229 39
o4-mini 84.7 0.35 0.6300 0.7915 0.8092 0.0156 31
GPT-4.1 71.2 0.39 0.4417 0.5745 0.5723 0.0142 29
GPT-4o 68.5 0.39 0.4303 0.5700 0.5610 0.0134 23
Open-source MLLMs
Qwen3-VL-235B-A22B-Thinking 75.1 0.38 0.4606 0.6021 0.5759 - 105
Qwen3-VL-235B-A22B-Instruct 71.6 0.43 0.4279 0.5703 0.6182 - 35
InternVL3.5-241B-A28B (thinking) 63.3 0.41 0.2856 0.3538 0.3866 29
Intern-S1 (241B, thinking) 57.0 0.51 0.0664* 0.0833* 0.2088* 39
Intern-S1-mini (8B, thinking) 56.6 0.53 0.0027* 0.0023* 0.0158* 18
Table 4: Ablation on the evaluator model. We use the subset in Tab. 3. Scoring points accuracy is the accuracy between model’s answers and human’s answers of the scoring points. Plausibility MAE is the average Mean Absolute Error of Spelling, Logical Consistency and Readability. Correlations are calculated similar to Tab. 3(b). *: The $p$-value is larger than 0.05, indicating the metric using this evaluator has no statistical significance.
We further study the selection of evaluator model by examining different MLLMs, including closed-source models (GPT-5, GPT-5-mini, o3, o4-mini, GPT-4.1, GPT-4o OpenAI (2025a; c)) and open-source models (Qwen3-VL Bai et al. (2025b), InternVL3.5 Wang et al. (2025a), Intern-S1, Intern-S1-mini Bai et al. (2025a)). On the same manually evaluated samples as in Tab. 3, we run each of the models as evaluator and calculate the accuracy of answering scoring points, MAE of visual plausibility, and three correlation metrics (Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$).
For open-source models, Qwen3-VL demonstrates relative strong performance, but still lags behind GPT-5. Other models have a larger gap between closed-source models, where some scoring point accuracy is close to random guess (50%) and some correlations are even of no statistical significance, suggesting these evaluators are not applicable. This is because multidisciplinary images are still difficult for current multimodal understanding models Yue et al. (2024b). Therefore, it is necessary to use advanced MLLMs to evaluate the generated images for higher precision and robustness.
4.5 Qualitative Analysis
Examples of generated images are shown in Fig. 7. Taking GPT-Image-1 as an example, common error types include lack of knowledge (e.g. keys in the Circle of Semitones), incorrect reasoning (e.g. failing to infer the coordinate of the intersection point), and graph-drawing errors of logical inconsistency (e.g. the point (1,0)(-1,0) appears to be on the y-axis), highlighting the necessity to integrate multimodal understanding, reasoning and generation. Some models show other graph-drawing errors of spelling (e.g. “Transpostion” for Qwen-Image) or readability (e.g. occluded “Music Theory” for HiDream-I1-Full), which is a key capability in multidisciplinary text-to-image exams.
These examples align with our quantitative results: while top closed-source models can partially integrate basic knowledge into generation, they still struggle with precise, domain-aware execution; open-source models, on the contrary, often fail at basic aspects like generating visual plausible images. This underscores the need for future models to prioritize multidisciplinary knowledge integration, rigorous reasoning, and fine-grained visual coherence. Such failures also demonstrate the value of GenExam ’s fine-grained scoring criteria, as they capture flaws that broad metrics like photorealism or general aesthetics would miss. More qualitative results are provided in Appendix G.
@@ 5 Conclusion
We presented GenExam, the first benchmark for multidisciplinary text-to-image exams, designed to test whether models can integrate understanding, reasoning and generation to truly solve graph-drawing problems. GenExam offers diverse and challenging prompts with ground truth images and fine-grained scoring points, reflecting the rigor of real exams. Our experiments demonstrate that even the strongest models struggle to achieve substantial performance, highlighting the difficulty of the benchmark. We hope GenExam will guide the advancement of generative models toward expert-level intelligence in a multidisciplinary direction.
Ethics Statement
All images in our benchmark are constructed from publicly available resources and existing datasets. We will release the benchmark with clear usage guidelines to prevent misuse for generating misleading content. We have checked for potential biases across all the subjects, aiming for balanced knowledge coverage without favoring specific demographics or disciplines.
@@ Appendix A More Statistics
Appendix B Automatic Data Filtering
To obtain a high-quality dataset, we employ an automatic filtering process on the collected images. We first design heuristic rules to remove images with duplicated content, low resolution, watermarks or non-English text. Then, we leverage GPT-5 to rate each image in terms of its text richness, image domain, image complexity, and subject knowledge density. We set a threshold for each of these aspects to filter out images that do not meet the requirements of our task, including those with:
@@ Appendix C Additional Experiments
In Tab. 5, we show the results on GenExam-Mini, which are consistent with the full dataset. GenExam-Mini can be considered as a valid subset for efficient and affordable validation.
@@
Results categorized by difficulty levels are presented in Tab. 7. We observe that models tend to have lower performance on samples with higher subject knowledge difficulty, demonstrating the challenge of integrate profound multidisciplinary knowledge into generation.
The results on each level-2 taxonomy is provided in Tab. 8. To highlight the difference between models, and since some models can only obtain a 0% strict score, we only report relaxed scores.
@@
Gemini-2.5-Flash-Image 4.3 58.7
Seedream 3.0 0.4 23.4
FLUX.1 Kontext max 0.0 31.6
@@
Table 5: Strict and relaxed scores on GenExam-Mini.
Model
Chemical
Structures
Diagrams
Geometric
Shapes
Maps
Plots &
Charts
Sheet
Music
Trees &
Graphs
Other
@@
Seedream 4.0 4.8 38.6 10.8 59.3 1.6 40.3 5.4 54.1 6.5 53.0 0.0 35.2 2.8 48.0 6.9 54.2
Gemini-2.5-Flash-Image 4.8 37.1 5.6 63.4 1.6 46.3 3.6 61.2 2.2 53.1 1.7 37.5 2.8 42.9 3.4 59.6
@@
Model Easy Medium Hard
Strict Relaxed Strict Relaxed Strict Relaxed
Closed-source Models
GPT-Image-1 18.6 68.1 12.8 62.2 9.9 58.5
Seedream 4.0 11.2 57.8 5.5 50.3 7.9 53.2
Gemini-2.5-Flash-Image 5.4 58.4 3.2 55.5 4.7 55.4
Imagen-4-Ultra 9.9 58.6 7.1 50.7 7.1 51.8
Seedream 3.0 0.4 31.3 0.3 21.8 0.0 19.0
FLUX.1 Kontext max 0.0 37.3 0.0 26.1 0.0 23.8
Open-source T2I Models
Qwen-Image 0.4 34.6 0.3 25.4 0.0 21.3
HiDream-I1-Full 0.0 29.7 0.0 19.8 0.0 17.1
FLUX.1 dev 0.0 25.2 0.0 15.4 0.0 12.3
FLUX.1 Krea 0.0 24.9 0.0 13.0 0.0 12.1
Stable Diffusion 3.5 Large 0.0 25.0 0.0 14.9 0.0 11.2
Open-source Unified MLLMs
BAGEL (thinking) 0.0 21.1 0.0 12.8 0.0 7.7
BAGEL 0.0 17.9 0.0 11.5 0.0 7.1
Show-o2-7B 0.0 18.1 0.0 10.2 0.0 7.9
Show-o2-1.5B-HQ 0.0 17.1 0.0 8.7 0.0 6.9
BLIP3o-NEXT-GRPO-Text-3B 0.0 20.2 0.0 12.3 0.0 8.1
BLIP3o-8B 0.0 10.6 0.0 5.8 0.0 5.1
Janus-Pro 0.0 13.6 0.0 9.9 0.0 6.5
Emu3 0.0 14.2 0.0 7.5 0.0 6.0
Table 7: Results categorized by subject knowledge difficulty.
@@
Seedream 4.0
Gemini-2.5-Flash-Image

### v3

@@ 3.2 Evaluation Protocol
Semantic Correctness Spelling Logical Consistency Readability
@@
Relaxed Score By GPT-5 0.6746 <<0.0001 0.8217 <<0.0001 0.8444 <<0.0001
Relaxed Score By Gemini-3-Flash 0.6608 <<0.0001 0.8200 <<0.0001 0.8261 <<0.0001
@@
Table 3: Mean absolute error (MAE) and correlations between human preferences and automatic evaluation.
4 Experiments
In this section, we evaluate the performance of representative methods on our GenExam. We also validate the alignment between MLLM-as-a-judge and human evaluation, and offer analysis on the failure cases of existing methods. We use GPT-5 (OpenAI, 2025a) as the evaluator model and set the reasoning effort as “low”. More experiments are provided in Appendix C, including results categorized by subject knowledge difficulty and level-2 taxonomy, detailed scores on each dimension, and ablations on the evaluators.
4.1 Baselines
We select 17 representative models for comparison, including closed-source proprietary models, open-source T2I models and unified MLLMs. We use the default T2I configuration for each of the models for inference.
Closed-source Models: Nano Banana Pro (Google, 2025b), GPT-Image-1.5 (OpenAI, 2025c), GPT-Image-1 (OpenAI, 2025b), Seedream 4.5 (Seed, 2025), FLUX.2 max (Labs, 2025), and Imagen-4-Ultra (Deepmind, 2025). These proprietary models represent the state-of-the-art methods in text-to-image generation.
Open-source T2I Models: Qwen-Image-2512 (Wu et al., 2025), FLUX.2 dev (Labs, 2025), HiDream-I1-Full (Cai et al., 2025), and Stable Diffusion 3.5 Large (Rombach et al., 2022). These models are dedicated to the T2I task.
Open-source Unified MLLMs: Show-o2-7B, Show-o2-1.5B-HQ (Xie et al., 2025), BAGEL (Deng et al., 2025) (thinking & non-thinking), Janus-Pro (Chen et al., 2025c), Emu3 (Wang et al., 2024a), BLIP3o-8B (Chen et al., 2025a), and BLIP3o-NEXT-GRPO-Text-3B (Chen et al., 2025b). These models employ a unified architecture for T2I and multimodal understanding.
4.2 Main Results
Results on GenExam are provided in Tab. 2. The strict scores demonstrate the challenge of our benchmark. While Nano Banana Pro and GPT-Image-1.5 show relatively strong performance of 72.7% and 43.2%, other closed-source methods like Seedream 4.5 and FLUX.2 max can only achieve strict scores less than 15%. For open-source models, even the latest Qwen-Image-2512 and FLUX.2 dev fail in almost all cases, with strict scores less than 5%. This shows a huge gap between open-source and closed-source models in multidisciplinary text-to-image exams, suggesting potential direction for future improvement.
The differences between the models are highlighted with relaxed scores. The performance gap between open-source and closed-source models still exists. FLUX.2 dev and Qwen-Image-2512 achieve the highest relaxed scores of 42.9 and 35.8 among open-source methods. However, current unified MLLMs like BAGEL and Show-o2 still lag behind dedicated T2I models, suggesting a significant challenge for unified MLLMs to utilize its multidisciplinary knowledge and reasoning ability into image generation.
@@ 4.3 Human Alignment
We conduct a human alignment experiment to examine the validity of MLLM-as-a-judge, following the paradigms in  (Zhao et al., 2025; Sun et al., 2025). Five experts are asked to score 250 randomly sampled images generated by GPT-Image-1.
@@
They first provide an overall rating between 1 and 10, and then follow the automatic evaluation protocol to annotate scoring points, spelling, logical consistency and readability. The average scores among all experts are used to calculate the mean absolute error (MAE) and correlations between MLLM-judge predicted scores and human scores. In Tab. 3(a), the results show a low MAE across the four dimensions, indicating high precision of our evaluation.
@@
We then use the overall rating to calculate correlations (Kendall’s $\tau$, Spearman’s $\rho$ and Pearson’s $r$) between human scores and four automatic metrics: our relaxed score, our semantic correctness, VQA score (Lin et al., 2024), and CLIP score (Hessel et al., 2021). In Tab. 3(b), our relaxed score evaluated by GPT-5 achieves high correlations with human judgments with $p<0.05$, slightly better than using solely semantic correctness without visual plausibility, while CLIP score and VQA score fail to capture the correctness of multidisciplinary images. When using Gemini-3-Flash (Google, 2025a) as the evaluator for relaxed scores, the results still show strong human alignment.
In Appendix Tab. 4, we further examine other closed-source and open-source evaluators. In Appendix Tab. 5, we extend to generated images of five other models to confirm strong correlations with human ratings across all models, and provide standard deviations to demonstrate the stability of the evaluation.
4.4 Error Analysis
To gain deeper insights into the strengths and limitations of the models, we provide a comparison of representative models on each evaluation dimension in Fig. 5. Specifically, Nano Banana Pro and GPT-Image-1.5 surpass other models across all four dimensions, notably with near-perfect spelling and readability, and Nano Banana Pro shows even stronger semantic correctness and logic consistency than GPT-Image-1.5. Open-source models Qwen-Image-2512 and FLUX.2 dev lag behind closed-source models across all dimensions, especially on spelling and logic consistency, suggesting that open-source models should first focus on basic drawing before tackling high-level semantics and reasoning.
Examples of generated images are shown in Fig. 6. As shown, state-of-the-art closed-source models such as Nano Banana Pro can generate relatively high-quality images that adheres to prompt specifications, despite some minor flaws. In contrast, open-source models often struggle with fundamental rendering skills, frequently generating incoherent figures characterized by legibility errors (e.g., the blurred label “C” in Qwen-Image-2512 in History), logical inconsistencies (e.g., redundant axes in Qwen-Image-2512 Math), and spelling inaccuracies (e.g., ”Ea?ies ?e’ws” in FLUX.2 dev in History). These visual defects not only diminish visual plausibility scores but also impede semantic evaluation.
It is worth noting that the inferior performance of open-source models often results not from a lack of necessary disciplinary knowledge, but from an inability to accurately translate that knowledge into readable visual figures. For instance, in the result of FLUX.2 dev in history, the model correctly identifies the regions of Egypt, Iran, India, and China (although with excessive highlighting) but fails to place the appropriate graphics within those regions. This discrepancy suggests that the bottleneck lies not in the lack of knowledge, but in the failure to execute visual reasoning based on that knowledge. Consequently, enhancing the ability to transform internal knowledge into visual outputs and improving the alignment between conceptual understanding and generative execution could improve open-source models to a new stage of performance. These specific failures further underscore the value of GenExam’s fine-grained scoring criteria, which effectively capture nuanced flaws that broad metrics such as photorealism or general aesthetics would typically overlook. Additional qualitative results are provided in Appendix F.
@@ 5 Conclusion
We presented GenExam, the first benchmark for multidisciplinary text-to-image exams, designed to test whether models can integrate understanding, reasoning and generation to truly solve graph-drawing problems. GenExam offers diverse and challenging prompts with ground truth images and fine-grained scoring points, reflecting the rigor of real exams. Our comprehensive experiments on 17 existing models demonstrate the difficulty of the benchmark and the gap between closed-source and open-source models. We hope GenExam will guide the advancement of generative models toward expert-level intelligence in a multidisciplinary direction.
@@ Appendix A More Statistics
Appendix B Data Curation Pipeline
Data Collection. The data curation pipeline is shown in Fig. 10. To construct a general benchmark for multidisciplinary text-to-image exam, we first consider possible subjects and domains of multidisciplinary images and curate a four-level taxonomy assisted by GPT-5 (OpenAI, 2025a) and through manual check. We then use these taxonomy to generate keywords related to specific subject knowledge, and collect web images from exams and textbooks based on Google search. Images from existing subject-knowledge-related datasets are also collected: MMMU (Yue et al., 2024a), ScienceQA (Lu et al., 2022), TextbookQA (Kembhavi et al., 2017), MAVIS (Zhang et al., 2024), ChEBI (Hastings et al., 2016), HLE (Phan et al., 2025) and R-Bench (Guo et al., 2025). We obtain 40K images from the data sources with their original text descriptions (e.g. webpage content for web images, lecture notes in TextbookQA, or question and answers in MMMU).
Automatic Filtering. To obtain a high-quality dataset, we employ an automatic filtering process on the collected images. We first design heuristic rules to remove images with duplicated content, low resolution, watermarks or non-English text. Then, we leverage GPT-5 to rate each image in terms of its text richness, image domain, image complexity, and subject knowledge density. We set a threshold for each of these aspects to filter out images that do not meet the requirements of our task, including those with:
@@ Appendix B Data Curation Pipeline
We obtain 6.5K images from automatic filtering for annotating.
Annotating. We employ GPT-5 to generate the initial version of prompts and scoring points. Prompts that require subject knowledge and reasoning are generated, similar to graph-drawing questions in exams. The original text description from the data source is used for reference to improve the precision, bu should be treated carefully since they may contain irrelevant information. Subsequently, scoring points (questions and scores) are generated based on the prompt. We enforce through prompt engineering that the questions cannot be too general (e.g. “Does the image show the correct structure of benzene?”) or too complex (can be split into multiple sub-questions). Scores are based on the difficulty of the questions, where basic questions should have low scores and harder ones have high scores. We provide some few-shot examples to enhance the labeling process. The complete prompts are provided in Appendix E.2.
@@ Appendix C Additional Experiments
C.1 Ablation on Evaluator Models
Evaluator Model Ground Truth Input Scoring Points Accuracy $\uparrow$ Plausibility MAE $\downarrow$ Kendall $\tau\uparrow$ Spearman $\rho\uparrow$ Pearson $r\uparrow$ API Cost (\$/img) Time (s/img)
Closed-source MLLMs
Gemini-3-Flash Img 86.4 - 0.6608 0.8300 0.8261 0.063 27
GPT-5 + low reasoning effort Img 88.5 0.20 0.6746 0.8217 0.8444 0.0242 72
GPT-5 + low reasoning effort - 86.2 0.23 0.6141 0.7753 0.7976 0.0231 59
GPT-5 + low reasoning effort Text 87.8 0.21 0.6559 0.8103 0.8336 0.0234 63
GPT-5 + low reasoning effort Img, Text 88.5 0.20 0.6693 0.8345 0.8427 0.0250 75
GPT-5 + minimal reasoning effort Img 84.2 0.28 0.6066 0.7521 0.7860 0.0144 35
GPT-5 + medium reasoning effort Img 86.7 0.21 0.6617 0.8224 0.8358 0.0389 101
GPT-5 + high reasoning effort Img 87.6 0.20 0.6647 0.8190 0.8304 0.0654 186
GPT-5-mini + minimal reasoning effort Img 79.0 0.35 0.5446 0.6996 0.7185 0.0039 11
GPT-5-mini + low reasoning effort Img 84.3 0.29 0.6002 0.7729 0.7806 0.0052 19
GPT-5-mini + medium reasoning effort Img 84.2 0.25 0.6313 0.8025 0.8165 0.0079 36
GPT-5-mini + high reasoning effort Img 83.4 0.26 0.5906 0.7611 0.7811 0.0196 119
o3 Img 82.0 0.23 0.6082 0.7706 0.7895 0.0229 39
o4-mini Img 84.7 0.35 0.6300 0.7915 0.8092 0.0156 31
GPT-4.1 Img 71.2 0.39 0.4417 0.5745 0.5723 0.0142 29
GPT-4o Img 68.5 0.39 0.4303 0.5700 0.5610 0.0134 23
Open-source MLLMs
Qwen3-VL-235B-A22B-Thinking Img 75.1 0.38 0.4606 0.6021 0.5759 - 105
Qwen3-VL-235B-A22B-Instruct Img 71.6 0.43 0.4279 0.5703 0.6182 - 35
InternVL3.5-241B-A28B (thinking) Img 63.3 0.41 0.2856 0.3538 0.3866 29
Intern-S1 (241B, thinking) Img 57.0 0.51 0.0664* 0.0833* 0.2088* 39
Intern-S1-mini (8B, thinking) Img 56.6 0.53 0.0027* 0.0023* 0.0158* 18
Table 4: Ablation on the evaluator model. Settings are identical to Tab. 3. *: The $p$-value is larger than 0.05, indicating the metric using this evaluator has no statistical significance.
We further study the selection of evaluator model by examining different MLLMs, including closed-source models (Gemini-3-Flash (Google, 2025a), GPT-5, GPT-5-mini, o3, o4-mini, GPT-4.1, GPT-4o (OpenAI, 2025a, d)) and open-source models (Qwen3-VL (Bai et al., 2025b), InternVL3.5 (Wang et al., 2025), Intern-S1, Intern-S1-mini (Bai et al., 2025a)). On the same manually evaluated samples as in Tab. 3, we run each of the models as evaluator and calculate the accuracy of answering scoring points, MAE of visual plausibility, and three correlation metrics (Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$).
For open-source models, Qwen3-VL demonstrates relative strong performance, but still lags behind GPT-5. Other models have a larger gap between closed-source models, where some scoring point accuracy is close to random guess (50%) and some correlations are even of no statistical significance, suggesting these evaluators are not applicable. This is because multidisciplinary images are still difficult for current multimodal understanding models (Yue et al., 2024b). Therefore, it is necessary to use advanced MLLMs to evaluate the generated images for higher precision and robustness.
We observe that using ground truth images as reference helps to improve the evaluation precision. We also test using captions of the ground truth images generated by GPT-5 as the ground truth, and observe that text-only ground truth is slightly worse than image-only ground truth. This is because ground truth images have better spatial/detail fidelity than textual descriptions (which often omit critical spatial details) and lower ambiguity. Using both images and textual descriptions shows no significant improvement than image-only, as images already contain sufficient detailed information and incorporating text becomes unnecessary.
C.2 Detailed Comparison of Existing Methods
In Tab. 5, we extend the human alignment experiment to 500 additional samples covering five other models. The results show that relaxed scores maintain strong correlations with human ratings across all models, confirming the metric’s generalizability. The standard deviation (std) values demonstrate high inter-rater reliability.
In Tab. 6 (a), we show the results on GenExam-Mini. The performance of the models is consistent with the results from the full dataset, which can be considered as a valid subset for efficient and affordable validation.
Results categorized by subject knowlege difficulty are presented in Tab. 6 (b). We observe that models tend to have lower performance on samples with higher subject knowledge difficulty, demonstrating the challenge of integrate profound multidisciplinary knowledge into generation.
@@ C.2 Detailed Comparison of Existing Methods
The results on each level-2 taxonomy are provided in Tab. 8. To highlight the difference between models, and since some models can only obtain a 0% strict score, we only report relaxed scores.
@@
Model Semantic Spelling Logic Readability Kendall $\tau$ Spearman $\rho$ Pearson $r$
MAE std MAE std MAE std MAE std
Nano Banana Pro 0.068 0.0142 0.071 0.0448 0.158 0.0573 0.118 0.0451 0.7246 0.8621 0.8893
GPT-1.5-Image 0.082 0.0169 0.086 0.0537 0.191 0.0692 0.143 0.0548 0.7013 0.8416 0.8679
Seedream 4.5 0.108 0.0224 0.113 0.0701 0.261 0.0946 0.195 0.0759 0.6672 0.8129 0.8248
FLUX.2 dev 0.126 0.0258 0.132 0.0817 0.301 0.1085 0.225 0.0853 0.6461 0.7928 0.7986
Qwen-Image-2512 0.134 0.0273 0.140 0.0864 0.317 0.1149 0.238 0.0901 0.6319 0.7784 0.7816
Table 5: Human alignment on images generated by more models.
@@
(a) GenExam-Mini
Model Easy Medium Hard
Strict Relaxed Strict Relaxed Strict Relaxed
Closed-source Models
Nano Banana Pro 73.1 93.8 71.8 92.8 66.8 92.7
GPT-Image-1.5 50.8 85.6 41.4 79.1 38.3 81.3
GPT-Image-1 18.6 68.1 12.8 62.2 9.9 58.5
Seedream 4.5 12.8 59.9 9.8 56.9 14.5 61.9
FLUX.2 max 14.5 66.0 7.7 61.0 5.8 59.3
Imagen-4-Ultra 9.9 58.6 7.1 50.7 7.1 51.8
Open-source T2I Models
FLUX.2 dev 2.9 48.9 2.1 40.4 2.4 39.9
Qwen-Image 0.4 34.6 0.3 25.4 0.0 21.3
HiDream-I1-Full 0.0 29.7 0.0 19.8 0.0 17.1
Stable Diffusion 3.5 Large 0.0 25.0 0.0 14.9 0.0 11.2
Open-source Unified MLLMs
BAGEL (thinking) 0.0 21.1 0.0 12.8 0.0 7.7
BAGEL 0.0 17.9 0.0 11.5 0.0 7.1
Show-o2-7B 0.0 18.1 0.0 10.2 0.0 7.9
Show-o2-1.5B-HQ 0.0 17.1 0.0 8.7 0.0 6.9
BLIP3o-NEXT-GRPO-Text-3B 0.0 20.2 0.0 12.3 0.0 8.1
BLIP3o-8B 0.0 10.6 0.0 5.8 0.0 5.1
Janus-Pro 0.0 13.6 0.0 9.9 0.0 6.5
Emu3 0.0 14.2 0.0 7.5 0.0 6.0
(b) Subject Knowledge Difficulty
Table 6: Results on GenExam-Mini and on different subject knowledge difficulty levels.
Model Chemical Structures Diagrams Geometric Shapes Maps Plots & Charts Sheet Music Trees & Graphs Other
@@
Seedream 4.5 4.8 44.8 16.4 66.2 1.6 42.4 0.0 54.2 15.2 63.8 3.3 38.1 11.1 52.0 10.3 60.1
FLUX.2 max 6.0 46.1 9.7 67.2 8.1 56.1 7.1 67.6 6.5 57.5 6.7 48.5 11.1 51.4 10.3 60.8
@@
Seedream 4.5
FLUX.2 max